You own the rights to two movies, Casablanca and Godzilla. Your accounting cost for these movies are $ 900 for Casablanca and $500 for Godzilla. You are planning on selling the rights to show the movies to two TV stations, Station A and Station B. You could them or selling them separately. You know that Station A serves mostly mature adults viewers and values Casablanca and Godzilla as follows (see table below). Station B serves mostly young adults and their willingness to pay for both movies is also revealed below.

Question: If arbitrage is not possible, what is the best way to sell these movies? Do you sell them as a bundle or single? What would be the best way of selling them to maximize your profit? (3 points)

Casablanca    Godzilla

Station A: $10,000    $1000

Station B: $2,000    $12,000
Answer #1

Let us consider the case of individual pricing.

In case of Casablanca, we have two possible prices $10,000 and $2000

If price is $2000, both stations will buy the product and profit=2*2000-2*900=$2200

If price is $10000, Station A will buy the product and profit=1*10000-1*900=$9100

We can see it is profit maximizing to sell Casablanca at a price of $10000

Let us take the case of Godzilla, we have two possible prices $12,000 and $1000

If price is $1000, both stations will buy the product and profit=2*1000-2*500=$1000

If price is $12000, Station B will buy the product and profit=1*12000-1*500=$11500

We can see it is profit maximizing to sell Godzilla at a price of $12000

Total Profit in case of individual pricing=9100+11500=$20600

Let us consider the case of bundle pricing. Let us consider the bundle of Casablanca and Godzilla. Cost of a bundle=900+500=$1400

Possible prices are $11000 and $14000

If price is $11000, both stations will buy the product and profit=2*11000-2*1400=$19200

If price is $14000, Station B will buy the product and profit=1*14000-1*1400=$12600

We can see it is profit maximizing to sell bundle at a price of $11000

We can see that it is more profitable to sell products as individual than selling it in bundle.